OpenStudy (anonymous):

A father invests certain amount of money at 6%. His son also invests one-third the amount his father invested, at 7%. If their combined income in one year is $250, how much did the son invest? a. $1000 b. $2000 c. $3000 d. $4000

OpenStudy (gw2011):

The answer is (a) $1000. Here is how I got the answer: Let x= father's investment 0.06x+(0.07)(1/3)(x)=250 0.06x+(0.07/3)(x)=250 Multiply by 3 to eliminate the denominator 0.18x+0.07x=750 0.25x=750 x=$3000 This is the father's investment (1/3)(3000)=$1000 This is the son's investment